OpenStudy (pinklion23):

Which answer shows this equation in standard form? 6 – 2(x – y) = –3x + 5 A.x + 2y = –1 B.x + 2y = 1 C.x – 2y = 1 D.–5x + 2y = –1

OpenStudy (pinklion23):

@paki

OpenStudy (pinklion23):

the first thing i did was take away the parenthesis

OpenStudy (anonymous):

first release the parantheses: 2(x-y)=2x-2y now -(2x-2y) equals 2y-2x So 6 - 2(x-y) will be equal to 6 + 2y - 2x

OpenStudy (pinklion23):

then i got 6-2x-2y=-3x+5

OpenStudy (pinklion23):

yeah i did that

OpenStudy (jhannybean):

standard form: Ax + By = C

OpenStudy (anonymous):

now here is your equation 6+2y-2x=-3x+5 add 2x to both sides: 6+2y-2x+2x=-3x+2x+5 Simplify: 6+2y=5-x subtract 6 from both sides: 6-6+2y=5-x-6 You will get: 2y=-x-1 Divide both sides by 2: y=(-x-1)/2 or y=-(x+1)/2

OpenStudy (pinklion23):

Thx jhanny i knew that :)

OpenStudy (anonymous):

oops, sorry then @Jhannybean and @pinklion23

OpenStudy (pinklion23):

?

OpenStudy (pinklion23):

jhannybean what do you think

OpenStudy (jhannybean):

\[6 – 2(x – y) = –3x + 5\]\[6-2x+2y= -3x+5\]\[(3-2)x +2y = 5-6\]Simplify.

OpenStudy (pinklion23):

ok

OpenStudy (anonymous):

ok, I will finish this equation. here is your equation: 6+2y-2x=-3x+5 add 3x to both sides: 6+2y-2x+3x=-3x+3x+5 Simplify: 6+2y+x=5 Subtract 6 from both sides: 6+2y+x-6=5-6 2y+x=-1

OpenStudy (pinklion23):

\[x+2y=-1\]

OpenStudy (anonymous):

yes. that's right

OpenStudy (pinklion23):

A is my answer then

OpenStudy (pinklion23):

hey great you guys gave medals to each other

OpenStudy (anonymous):

yes.

OpenStudy (pinklion23):

thx i have more though :0